# Jonathan Roberts1, Kai Han2, Samuel Albanie

### v2

@@ Abstract
[…]
@@ 1 Introduction
Figure 1: Overall performance on GRAB. Our benchmark proves challenging for frontier LMMs. The highest performing LMM attains an accuracy of just 21.7% on GRAB.
Figure 2: The GRaph Analysis Benchmark consists of 2170 synthetically generated graph analysis questions that prove challenging for frontier LMMs. The questions cover 23 graph properties organised into four core tasks: (i) Properties focuses on the analysis of features of individual functions and series; (ii) Functions and (iii) Series require computing the mean of properties across multiple functions and series; and (iv) Transforms involves determining the properties of a function after it has undergone a series of transforms.
@@
To this end, we introduce GRaph Analysis Benchmark (GRAB), consisting of 2170 questions. All figures in GRAB are generated synthetically using the Matplotlib library [13]. Although some realism might be lost when solely considering synthetically generated data, there are numerous benefits to taking this approach: (1) Difficulty – the complexity of questions can be controlled and tuned; (2) Noise – ground truth is determined automatically during curation, avoiding the need for potentially erroneous post hoc annotation; (3) Composition – we can directly select the characteristics of questions in our benchmark and arbitrarily control the scale; and, (4) Contamination – while the same figure style will almost certainly have been seen during the pretraining of most LMMs, it is unlikely that the exact figures and questions generated for our benchmark have.
Leveraging the control enabled by our synthetic generation, we optimise our benchmark for usefulness and longevity. Specifically, we include challenging questions that frontier LMMs struggle to answer (see Fig. 1) and we control the aesthetic parameters of each graph to ensure all relevant features are clear and questions are answerable. We restrict the size of our benchmark to 2170 questions, balancing streamlined evaluation and robustness.
We aim for a comprehensive evaluation of the abilities of frontier models to analyse figures by including a broad range of chart types and question formats. The tasks in GRAB involve determining key data properties, such as mean, variance, and interquartile range, as well as function properties like gradients, stationary points, and function parameters. Question complexity is increased by incorporating multiple functions and series (see Fig. 2) and increasing the required answer precision. We evaluate 20 closed source LMMs on our benchmark, which proves extremely challenging for the current generation of models, with the best performing attaining a score of 21.7%. While some models exhibit a modicum of success on the easiest question set, the vast majority of our benchmark is beyond the capabilities of all models. We conduct a detailed error analysis and set of ablations to gain insights into which tasks and categories were the most challenging, as well as investigations of how question format and precision impact performance.
We summarise the contributions of this work as follows: (i) We introduce GRAB (GRaph Analysis Benchmark), a challenging 2170 question benchmark for graph analysis, (ii) We comprehensively characterise the capabilities of 20 current generation frontier LMMs on GRAB, and (iii) We provide insights into model strengths and limitations through numerous ablations.
@@ 2.1 Figure Interpretation Benchmarks
Motivated by the last decade of progress in computer vision, a range of prior work has sought to evaluate figure the interpretation capabilities of leading models. One family of work has focused on synthetic data generation pipelines, profiting from the reliable ground truth and fine-grained control enabled by this approach. This has yielded figure binary classification tasks [15], captioning [7], and various structural understanding, retrieval and reasoning tasks with limited vocabularies [14]. Synthetic figure generation has also been combined with corpora of online statistics [6, 42, 25] and Kaggle datasets [46] to achieve greater figure diversity. A second line of work has sought to construct datasets by curating ‘real’ figures directly from online sources and parsing their metadata. For this purpose, figures have been sourced from research papers [41, 40] and online sources such as Our World in Data [23].
[…]
draws particular inspiration from the family of synthetic benchmarks, to which it also belongs. However, a key distinction between our tasks and prior benchmarks is GRAB’s significantly greater difficulty, as evidenced by our empirical results illustrating that even the strongest frontier LMMs struggle to make headway (Sec. 4).
@@ 2.2 LMM Evaluation and Benchmarks
Since the emergence of LMMs, numerous benchmarks have been created to evaluate and characterise their capabilities across many domains. Interpreting and answering questions of natural images is a key task, with benchmarks such as MMBench [19], SEEDBench [16], SEEDBench-2 [17], MME [9], MM-Vet [47], LVLM-eHub [45], ScienceQA [21] and MMMU [48] finding prominence. However, as outlined in [8], although widely used, some of these most popular benchmarks contain problematic questions, including those that do not require visual content to be answered. Unlike older datasets developed to cater to the supervised learning paradigm that typically required much larger scales, more recent LMM-focused benchmarks contain significantly smaller test sets of a few thousand questions or a few hundred (e.g. [35]).
Various benchmarks can be found in more niche domains, such as geographic and geospatial evaluation [39], hallucinations [12] and OCR [20]. A closely related work to GRAB is MathVista [22], which evaluates mathematical reasoning in visual contexts. MathVista is a metadataset (similar to [38]) consisting of a broad range of mostly multiple-choice questions from 31 multimodal datasets covering a diverse set of figure types and mathematical tasks ranging from table interpretation to puzzles. Though similar, the focus of MathVista on general mathematical reasoning, is distinct from ours, which is centered on the more narrow task of deriving and analysing graph properties. Moreover, our benchmark is considerably more challenging: GPT-4o scores >60% on MathVista [^1] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard compared to 13.6% on GRAB.
Statistic Number
Questions 2170
Tasks 4
Graph properties 23
Properties questions 660
Functions questions 710
Series questions 490
Transforms questions 310
10s precision questions 30
Integer precision questions 1575
1 d.p. precision questions 565
@@
Category Properties
Intercepts & Gradients x-intercept, y-intercept, gradient
Stationary Points stationary point(s) coordinates
Trigonometric amplitude, vertical shift, period
Functions function equations
Counting number of points, number of series
Correlation {Pearson’s, Spearman’s rank, Kendall rank} correlation coefficient
Area Bounded total area bounded, net area bounded
Measures of Spread mean, median, interquartile range, variance
Range & Extrema min/max values, domain length, range
Table 2: Categories and properties
Figure 3: GRAB category distribution
3 GRaph Analysis Benchmark
Our GRaph Analysis Benchmark consists of 2170 questions covering four core tasks and 23 graph properties. In the following section, we provide an overview of the benchmark, including descriptions of the categories and tasks, example questions and answers, and an overview of our synthetic curation process. High-level statistics can be found in Tab. 1.
3.1 Question format
All questions in our benchmark follow a single-answer format, in which each question asks for a specific property of the function(s) or data present in the graph along with required precision levels.
@@ 3.2 Categories
Questions in our GRAB benchmark cover 23 graph properties from which analytical insights can commonly be derived from data or functions in graphs and charts. We group these properties into the categories shown in Tab. 2. We consider this set of properties to cover most real use cases and require some measure of visual analytical reasoning. We refrain from including ‘easier’ OCR-centered questions – such as legend, title, axis-label interpretation – in our benchmark to keep the focus on visual analytical reasoning, and instead only include questions of this type as an ablation (see Appendix). Furthermore, as outlined in Sec. 2, numerous prior benchmarks have included this type of question.
@@ 3.3 Tasks
We structure GRAB into the following tasks covering different aspects of graph analysis. Example questions and answers for each task can be found in Fig. 4.
@@ 3.3.1 Properties
This initial task aims to provide a broad evaluation of the capability to understand and determine or derive properties of graphs, including both basic fundamental properties like number of series or more advanced properties requiring calculations, like total area bounded. Aside from number of series questions, the graphs involve just a single function or data series.
@@ 3.3.4 Transforms
Questions in the Transforms task includes a single function and require computing a graph property after performing a sequence of up to 10 transforms (rotations, translations, scales and reflections).
@@ 3.4 Curation
We leverage the Matplotlib [13] library to synthetically create the graphs used in our core benchmark and ablations studies. The following provides a brief overview of the methodology used to generate the figures and questions. For each task, we initially created a longlist of 250 questions per property. We then performed downsampling to select a final question set with uniformly distributed answers. This downselection aimed to increase the question diversity, reduced bias in the generation process and avoid an overabundance of questions with answers converging towards 0, especially for the Functions, Series and Transforms tasks.
We associate each graph property (Tab. 2) with a set of data generation functions and function parameter ranges. For categories such as Intercepts & Gradients or Functions, we sample values for the property of interest and then fit the generation parameters to produce functions/series that match the value. For the other categories, such as Correlation or Area Bounded, we randomly sample the generation parameters and then calculate the property value from the data. In each case, we ensure questions are unique. For each graph in the Properties task, aesthetic parameters (e.g. figure size, font size, and line style) are randomly sampled from a defined set of options. For the other tasks (Series, Functions and Transforms), we limit the variability of the graphs by configuring the appearance to a defined schema. As such, all graphs have the same x and y-limits, grid lines, tick labels, font and resolution (2251x2171). A single function type is used for each graph, irrespective of the number of series.
Most questions require answers to integer precision, though to increase difficulty, approximately 25% of questions require 1 decimal place precision. Questions in the Area Bounded and Functions categories are evaluated to either integer or nearest 10 precision.
Figure 4: Example questions for the four tasks in the GRAB benchmark.
@@ 4.2.1 Inference
We evaluate the closed-source models via the Vertex AI API [11] {Gemini, Claude}, OpenAI API [30] {GPT}, and Reka API [2] {Reka}. Following benchmarking conventions, we set model hyperparameters that result in deterministic and reproducible generation. This entails utilising a greedy search decoding strategy in which the most probable token is selected from the model vocabulary $V$ at each step, conditional on the preceding tokens i.e., $w_{n+1}=\arg\max_{w\in V}P(w|w_{1},w_{2},\ldots,w_{n})$. We achieve this by setting the $temperature$ and $topk$ parameters to zero and 1, respectively, and specifying random seeds. We run the open-source models using HuggingFace Transformers [44] and the OpenCompass toolkit [33].
@@ 4.3 Overall results
Task Properties Functions Series Transforms Overall

### v3

# GRAB: A Challenging GRaph Analysis Benchmark for Large Multimodal Models

@@ Abstract
Large multimodal models (LMMs) have exhibited proficiencies across many visual tasks. Although numerous well-known benchmarks exist to evaluate model performance, they increasingly have insufficient headroom. As such, there is a pressing need for a new generation of benchmarks challenging enough for the next generation of LMMs. One area that LMMs show potential is graph analysis, specifically, the tasks an analyst might typically perform when interpreting figures such as estimating the mean, intercepts or correlations of functions and data series. In this work, we introduce GRAB, a graph analysis benchmark, fit for current and future frontier LMMs. Our benchmark is predominantly synthetic, ensuring high-quality, noise-free questions. GRAB is comprised of 3284 questions, covering five tasks and 23 graph properties. We evaluate 20 LMMs on GRAB, finding it to be a challenging benchmark, with the highest performing model attaining a score of just 21.0%. Finally, we conduct various ablations to investigate where the models succeed and struggle. We release GRAB and a lightweight GRAB-Lite to encourage progress in this important, growing domain.
@@ 1 Introduction
Figure 1: Overall performance on GRAB. Our benchmark proves challenging for frontier LMMs. The highest performing LMM attains an accuracy of just 21.0% on GRAB.
Figure 2: The GRaph Analysis Benchmark consists of 3284 graph analysis questions that prove challenging for frontier LMMs. The questions cover 23 graph properties organised into five core tasks: (i) Properties focuses on the analysis of features of individual functions and series; (ii) Functions and (iii) Series require computing the mean of properties across multiple functions and series; (iv) Transforms involves determining the properties of a function after it has undergone a series of transforms; and, (v) Real encompasses question styles from the other tasks in more realistic formats, including sketched on whiteboard or paper, embedded in digital contexts or with added noise.
@@
To this end, we introduce GRaph Analysis Benchmark (GRAB), consisting of 3284 questions. Almost all figures in GRAB are generated synthetically using the Matplotlib library [22]. Although some realism might be lost when focusing on synthetically generated data, there are numerous benefits to taking this approach: (1) Difficulty – the complexity of questions can be controlled and tuned; (2) Noise – ground truth is determined automatically during curation, avoiding the need for potentially erroneous post hoc annotation; (3) Composition – we can directly select the characteristics of questions in our benchmark and arbitrarily control the scale; and, (4) Contamination – while the same figure style will almost certainly have been seen during the pretraining of most LMMs, it is unlikely that the exact figures and questions generated for our benchmark have. The inclusion of a 1114-question ‘Real’ task with figures that are hand-drawn, have added noise or are placed in digital screenshots increases the realism of the benchmark by introducing artefacts, distractors and camera effects.
Leveraging the control enabled by our synthetic generation, we optimise GRAB for usefulness and longevity. Specifically, we include challenging questions that frontier LMMs struggle to answer (see Fig. 1) and we control the aesthetic parameters of each graph to ensure all relevant features are clear and questions are answerable. To balance streamlined evaluation and robustness, we include 3284 questions in GRAB. We also provide a smaller 500-question version, called GRAB-Lite, as a lighter-weight evaluation.
We aim for a comprehensive evaluation of the abilities of frontier models to analyse figures by including a broad range of chart types and question formats. The tasks in GRAB involve determining key data properties, such as mean, variance, and interquartile range, as well as function properties like gradients, stationary points, and function parameters. Question complexity is increased by incorporating multiple functions and series (see Fig. 2) and increasing the required answer precision. We evaluate 20 LMMs on our benchmark, which proves extremely challenging for the current generation of models, with the best performing attaining a score of 21.0%. While some models exhibit a modicum of success on the easiest question set, the vast majority of our benchmark is beyond the capabilities of all models. We conduct a detailed error analysis and set of ablations to gain insights into which tasks and categories were the most challenging, as well as investigations of how question format and plotting library impact performance.
We summarise the major contributions of this work as follows: (i) We introduce GRAB, a challenging 3284 question benchmark for graph analysis, along with a smaller 500-question GRAB-Lite version (ii) We comprehensively characterise the capabilities of 20 current generation frontier LMMs on GRAB, and (iii) We provide insights into model strengths and limitations through numerous ablations.
@@ 2.1 Figure interpretation benchmarks
Motivated by the last decade of progress in computer vision, a range of prior work has sought to evaluate figure interpretation capabilities of leading models. One family of work has focused on synthetic data generation pipelines, profiting from the reliable ground truth and fine-grained control enabled by this approach. This has yielded figure binary classification tasks [24], captioning [10], and various structural understanding, retrieval and reasoning tasks with limited vocabularies [23]. Synthetic figure generation has also been combined with corpora of online statistics [9, 54, 34] and Kaggle datasets [61] to achieve greater figure diversity. A second line of work has sought to construct datasets by curating ‘real’ figures directly from online sources and parsing their metadata. For this purpose, figures have been sourced from research papers [53, 52] and online sources such as Our World in Data [32]. GRAB draws particular inspiration from the family of synthetic benchmarks, to which it also belongs. However, a key distinction between our tasks and prior works is GRAB’s significantly greater difficulty, as evidenced by our empirical results illustrating that even the strongest frontier LMMs struggle to make headway (§4).
@@ 2.2 LMM evaluation and benchmarks
Since the emergence of LMMs, numerous benchmarks have been created to evaluate and characterise their capabilities across many domains. Interpreting and answering questions of natural images is a key task, with benchmarks such as MMBench [28], SEEDBench [25], SEEDBench-2 [26], MME [12], MM-Vet [62], LVLM-eHub [60], ScienceQA [30] and MMMU [63] finding prominence. However, as outlined in [11], although widely used, some of these most popular benchmarks contain problematic questions, including those that do not require visual content to be answered. Unlike older datasets developed to cater to the supervised learning paradigm that typically required much larger scales, more recent LMM-focused benchmarks contain significantly smaller test sets of a few thousand questions or a few hundred (e.g., [47]). Various benchmarks can be found in more niche domains, such as geographic and geospatial evaluation [51], hallucinations [21] and OCR [29]. A closely related work to GRAB is MathVista [31], which evaluates mathematical reasoning in visual contexts. MathVista is a metadataset (similar to [50]) consisting of a broad range of mostly multiple-choice questions from 31 multimodal datasets covering a diverse set of figure types and mathematical tasks ranging from table interpretation to puzzles. Though similar, the focus of MathVista on general mathematical reasoning, is distinct from ours, which is centered on the more narrow task of deriving and analysing graph properties. Moreover, our benchmark is considerably more challenging: GPT-4o scores >60% on MathVista [^1] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard compared to 14.9% on GRAB.
3 GRaph Analysis Benchmark
Statistic Number Questions 3284 Tasks 5 Graph properties 23 Properties questions 660 Functions questions 710 Series questions 490 Transforms questions 310 Real questions 1114 10s precision questions 46 Integer precision questions 2450 1 d.p. precision questions 787
Category Properties Intercepts & Gradients x-intercept, y-intercept, gradient Stationary Points stationary point(s) coordinates Trigonometric amplitude, vertical shift, period Functions function equations Counting number of points, number of series Correlation {Pearson’s, Spearman’s rank, Kendall rank} correlation coefficient Area Bounded total area bounded, net area bounded Measures of Spread mean, median, interquartile range, variance Range & Extrema min/max values, domain length, range
@@ 3 GRaph Analysis Benchmark
Table 2: GRAB categories and graph properties
Figure 3: GRAB categories
Our GRaph Analysis Benchmark consists of 3284 questions covering five core tasks and 23 graph properties. In the following section, we provide an overview of the benchmark, including descriptions of the categories and tasks, example questions and answers, and an overview of our synthetic curation process. High-level statistics can be found in Tab. 1.
@@ 3.1 Categories
Questions in our GRAB benchmark cover 23 graph properties from which analytical insights can commonly be derived from data or functions in graphs and charts. We group these properties into the categories shown in Tab. 2. We consider this set of properties to cover most real use cases and require some measure of visual analytical reasoning. We refrain from including ‘easier’ OCR-centered questions – such as legend, title, axis-label interpretation – in our benchmark to keep the focus on visual analytical reasoning, and instead only include questions of this type as an ablation. Furthermore, as outlined in §2, numerous prior benchmarks have included this type of question.
@@ 3.2 Tasks
We structure GRAB into the following tasks covering different aspects of graph analysis. Example questions and answers for each task can be found in Figs. 2 and  4.
@@ 3.2.1 Properties
This initial task aims to provide a broad evaluation of the capability to understand and derive properties of graphs, including both basic fundamental properties like number of series or more advanced properties requiring calculations, like total area bounded. Aside from number of series questions, the graphs involve just a single function or data series.
@@ 3.2.4 Transforms
Questions in the Transforms task include a single function and require computing a graph property after performing a sequence of up to 10 transforms (rotations, translations, scales and reflections).
3.2.5 Real
Organised into four splits, the Real task modifies the other tasks with increased realism. This task includes Properties-style questions with figures hand-drawn on (1) Paper and (2) Whiteboards, and questions styled on all the other tasks with figures placed in (3) Screenshots (e.g., in emails, slides, documents and video calls) or with arbitrary (4) Noise added to the figures (e.g., artefacts, flips and blurs).
@@ 3.3 Curation
We leverage the Matplotlib [22] and Seaborn [56] libraries to synthetically create the majority of the graphs in GRAB and hand-draw only a small subset of the Real task. For each task, we initially created a longlist of 250 questions per graph property. We then performed downsampling to select a final question set with uniformly distributed answers. This downselection aimed to increase question diversity, reduce bias in the generation process and avoid an overabundance of questions with answers converging towards 0. We associate each graph property (Tab. 2) with a set of data generation functions and function parameter ranges. For categories such as Intercepts & Gradients or Functions, we sample values for the property of interest and then fit the generation parameters to produce functions/series that match the value. For the other categories, such as Correlation or Area Bounded, we randomly sample the generation parameters and calculate the property value from the data. In each case, we ensure questions are unique. For each graph in the Properties task, aesthetic parameters (e.g. figure/font size and line style) are randomly sampled from a defined set of options. For the Series, Functions and Transforms tasks, we limit the variability of the graphs by configuring the appearance to a defined schema: all graphs have the same x and y-limits, grid lines, tick labels, font and resolution (2251x2171). A single function type is used for each question, irrespective of the number of series. Synthetically generated questions from the Real task leveraged the generation procedure from the other tasks to create new figures. These figures were then embedded in digital screenshots (Screenshots split) or manipulated with visual noise (Noise split). Figures in the Paper and Whiteboard splits were hand-drawn using a range of functions and graph styles and subsequently photographed.
For each task, we include sufficient questions to represent different combinations of graph properties and question complexity. We aim for an approximate balance per graph property, resulting in the category distribution shown in Fig. 3. Most questions require answers to integer precision, though some categories (e.g., Area Bounded) are evaluated to nearest 10 precision. To increase the difficulty, $\sim$25% of questions require 1 decimal place precision.
@@ 4 Experiments
Figure 4: Example questions for the five tasks in the GRAB benchmark. All questions include synthetically rendered graphs apart from the Whiteboard and Paper splits of the Real task, which are hand-drawn and photographed.
@@ 4.2.1 Inference
We evaluate the closed-source models via the Vertex AI API [14] {Gemini, Claude}, OpenAI API [40] {GPT}, and Reka API [2] {Reka}. Following benchmarking conventions, we set model hyperparameters that result in deterministic and reproducible generation. This entails utilising a greedy search decoding strategy in which the most probable token is selected from the model vocabulary $V$ at each step, conditional on the preceding tokens i.e., $w_{n+1}=\arg\max_{w\in V}P(w|w_{1},w_{2},\ldots,w_{n})$. We achieve this by setting the temperaturetemperature and topktopk parameters to zero and 1, respectively, and specifying random seeds. We run the open-source models using HuggingFace Transformers [57] and the OpenCompass toolkit [45].
@@ 4.3 Overall results
Properties Functions Series Transforms Real Overall
Model (660) (710) (490) (310) (1114) (3284)
Closed-source LMMs
Claude 3 Haiku [3] 14.2 6.6 8.8 3.9 9.2 9.1
Claude 3 Sonnet [3] 15.3 8.6 4.5 4.8 12.4 10.3
Claude 3.5 Sonnet [4] 41.8 15.5 11.0 10.0 19.6 21.0
GPT-4 Turbo [36] 18.5 8.5 4.9 3.5 7.5 9.2
GPT-4o mini [38] 15.8 6.8 5.7 2.9 4.0 7.1
GPT-4o [37] 24.7 10.8 9.2 3.5 17.3 14.9
Gemini 1.0 Pro Vision [13] 20.2 5.8 6.9 6.1
Gemini 1.5 Flash [49] 28.5 11.5 8.4 9.0 17.1 16.1
Gemini 1.5 Pro [49] 34.2 11.4 13.3 6.5 20.3 18.8
Reka Edge [46] 11.8 8.7 11.6 1.9
Reka Flash [46] 13.2 10.1 6.3 3.9 10.0 9.5
Reka Core [46] 1.7 0.0 4.3 0.3 1.3 1.5
Open-source LMMs
CogVLM-Chat [55] 7.0 4.9 5.1 3.9 10.5 7.2
Qwen-VL-Chat [8] 10.2 6.6 5.1 2.9 4.6 6.1
OmniLMM-3b [44] 6.7 4.9 4.1 4.5 6.2 5.5
TransCore-M [48] 7.9 9.2 7.6 3.9 8.2 7.9
Yi-VL-6b [1] 5.6 8.6 7.1 4.2 9.7 7.7
Yi-VL-34b [1] 7.6 5.9 5.5 2.3 7.5 6.4
LLaVA-1.5 7b [27] 4.7 7.5 6.5 4.8 8.5 6.9
LLaVA-1.5 13b [27] 5.0 7.7 8.4 3.9 8.9 7.3
Table 3: Results on GRAB. The highest open and closed-source model scores for each task are highlighted and bold. unable to perform full evaluation due to model deprecation. Overall scores are weighted averages based on task size.
We present the results on GRAB in Tab. 3. A clear observation is that our benchmark proves challenging for frontier models, with low overall performances from all models (see also Fig. 1). The highest performing model, Claude 3.5 Sonnet, achieves a score of just 21.0%. As expected, the highest scores are attained on the Properties task, which includes questions that just require the derivation of a single graph property. Lower scores are achieved on the Functions and Series tasks, likely due to the added complexity of computing the mean of a given property over multiple functions/series. There is a macro-trend of slightly higher accuracy on the functions task compared to the series task, possibly due to the irregularity introduced by the noise in the series data points. Significantly lower accuracy is attained on the Transforms task. Reasonable scores are attained on the Real task, which includes questions from the other tasks, that approximately correspond to the average scores on the other tasks. This indicates that the Real questions are of a similar difficulty.
A contributing factor to these scores is the degree to which the models follow output format instructions. Given our exact string matching evaluation protocol, any deviation from the expected output is deemed incorrect. Despite potentially having superior mathematical reasoning abilities, models that struggle to precisely follow instructions or have proclivities to generate verbose answers suffer severe performance degradation. With the exception of Reka Core (which showed poor output format instruction following capabilities), all closed-source models outperformed the leading open-source model. However, the open-source models did consistently generate answers in the correct format. In this low-scoring regime where even the best models struggle, a comparison of model performance should be made with the caveat that all models can barely perform the tasks in the benchmark. We anticipate our benchmark will yield most benefit when evaluating the next generation of LMMs.
4.4 Real splits analysis
In Tab. 4, we report model scores on questions in each split in the Real task. The highest scores are attained on the Whiteboard and Paper splits, which include only less complex Properties-style questions. Notably, Claude 3.5 Sonnet struggles, scoring higher on the other splits. In general, model scores on both the Screenshots and Noise splits are comparable to the overall GRAB scores. This suggests the models are well-suited to dealing with real-world images, as the addition of visual distortions due to the noise artefacts or distractors and additional context in the screenshots does not significantly degrade performance. We also report the performance of two newer releases of the Gemini 1.5 Flash model (2.0 and 2.5), each showing subsequent improvements.
Accuracy (%)
Whiteboard Paper Screenshots Noise Overall